### v2

# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

Abstract
We consider edge modification problems towards block and strictly chordal graphs, where one is given an undirected graph $G=(V,E)$ and an integer $k\in\mathbb{N}$ and seeks to edit (add or delete) at most $k$ edges from $G$ to obtain a block graph or a strictly chordal graph. The completion and deletion variants of these problems are defined similarly by only allowing edge additions for the former and only edge deletions for the latter. Block graphs are a well-studied class of graphs and admit several characterizations, e.g. they are diamond-free chordal graphs. Strictly chordal graphs, also referred to as block duplicate graphs, are a natural generalization of block graphs where one can add true twins of cut-vertices. Strictly chordal graphs are exactly dart and gem-free chordal graphs. We prove the NP-completeness for most variants of these problems and provide:
  • $O(k^{2})$ vertex-kernels for Block Graph Editing and Block Graph Deletion
  • $O(k^{3})$ vertex-kernels for Strictly Chordal Completion and Strictly Chordal Deletion
  • an $O(k^{4})$ vertex-kernel for Strictly Chordal Editing
1 Introduction
Parameterized algorithms are among the most natural approaches to tackle NP-hard optimization problems [15]. In particular, they have been very successful in dealing with so-called edge modification problems on graphs: given as input an arbitrary graph $G=(V,E)$ and a parameter $k\in\mathbb{N}$, the goal is to transform $G$ into a graph with some specific properties (i.e., belonging to a specific graph class $\mathcal{G}$) by adding and/or deleting at most $k$ edges. Parameterized algorithms (also called FPT for fixed parameter tractable) aim at a time complexity of type $f(k)\cdot n^{O(1)}$, where $f$ is some computable function, hence the combinatorial explosion is restricted to parameter $k$.
When the target class $\mathcal{G}$ is characterized by a finite family of forbidden induced subgraphs, modification problems are FPT by a result of Cai [10]. Indeed, as long as the graph contains one of the forbidden subgraphs, one can try each possibility to correct this obstruction and branch by recursive calls. On each branch, the budget $k$ is strictly diminished, therefore the whole algorithm has a number of calls bounded by some function $f(k)$. The situation is more complicated when the target class $\mathcal{G}$ is characterized by an infinite family of forbidden induced subgraphs. Nonetheless, a large literature is devoted to edge modification problems towards chordal graphs (where we forbid all induced cycles with at least four vertices) as well as sub-classes of chordal graphs, typically obtained by requiring some fixed set of obstructions, besides the long cycles. Observe that, in this case, the situation remains relatively simple if we restrict ourselves to edge completion problems, where we are only allowed to add edges to the input graphs. Indeed, in this case, if a graph has an induced cycle of length longer than $k+3$, it cannot be made chordal by adding at most $k$ edges. Therefore we can use again the approach of Cai to deal with cycles of length at most $k+3$ and other obstructions, and either the algorithm finds a solution in $f(k)$ recursive calls, or we can conclude that we have a no-instance. The cases of edge deletion problems (where we are only allowed to remove edges) and edge editing problems (where we are allowed to both remove edges or add missing edges) are more complicated, since even long cycles can be eliminated by a single edge removal. Therefore more efforts and more sophisticated techniques were necessary in these situations, but several such problems turned out to be FPT [12, 16, 17]. The interested reader can refer to [13] for a broad and comprehensive survey on parameterized algorithms for edge modification problems.
We focus here on a sub-family of parameterized algorithms, namely on kernelization. The goal of kernelization is to provide a polynomial algorithm transforming any instance $(I,k)$ of the problem into an equivalent instance $(I^{\prime},k^{\prime})$ where $k^{\prime}$ is upper bounded by some function of $k$ (in our case we will simply have $k^{\prime}\leqslant k$), and the size of the new instance $I^{\prime}$ is upper bounded by some function $g(k)$. Hence the size of the reduced instance does not depend on the size of the original instance. Kernels are obtained throught a set of reduction rules. While kernelization is possible for all FPT problems (the two notions are actually equivalent), the interesting question is whether a given FPT problem admits polynomial kernels, where the size of the reduced instance is bounded by some polynomial in $k$. Note that, under some complexity assumptions, not all FPT problems admit polynomial kernels [6, 7, 8, 11, 24, 31].
In this paper we focus on Block Graph Editing, Block Graph Deletion and all three variants of modification problems towards striclty chordal graphs. Block graphs are a well-studied subclass of chordal graphs, that are diamond-free (Figure 1). They are moreover the graphs in which every biconnected component induces a clique. Notice that this observation allows for a simple polynomial-time algorithm for Block Graph Completion since one needs to turn every biconnected component of the input graph into a clique. Strictly chordal graphs are another subclass of chordal graphs, also known as block duplicate graphs [28, 29, 23]. They can be obtained from block graphs by repeatedly choosing some cut-vertex $u$ and adding a true twin $v$ of $u$, that is a vertex $v$ adjacent to $u$ and all neighbors of $u$ [23]. They can also be characterized as dart, gem-free chordal graphs (see Figure 1 and next section) and are thus ptolemaic (i.e. chordal distance-hereditary [9]). Strictly chordal graphs are also known to be a subclass of 4-leaf power graphs [29], and a super-class of 3-leaf power graphs [16].
Related work
Kernelization for chordal completion goes back to the ’90s and the seminal paper of Kaplan, Shamir and Tarjan [27]. In this work, the authors provide an $O(k^{3})$ vertex-kernel for Chordal Completion (also known as Minimum Fill-In) that was later improved to $O(k^{2})$ by a tighter analysis [35]. Since then, several authors addressed completion, deletion and/or editing problems towards sub-classes of chordal graphs, as 3-leaf power graphs [4], split and threshold graphs [25], proper interval graphs [5], trivially perfect graphs [3, 18, 21, 25, 19] or ptolemaic graphs [14]. All these classes have in common that they can be defined as chordal graphs, plus a constant number of obstructions. Several questions remain open, for example it is not known whether chordal deletion or chordal editing admit polynomial kernels [13]. If instead of modifying edges we want to remove vertices, the problem Block Graph Vertex Deletion has a kernel with $O(k^{4})$ vertices [1].
Our contribution
We first prove that most variants of the considered problems are NP-Complete, the only exception being the Block Graph Completion problem that admits a polynomial-time algorithm. Secondly, we give kernelization algorithms for all these problems, with different bounds. First of all, we begin by illustrating the techniques we will use on the Block Graph Editing problem and obtain an $O(k^{2})$ vertex-kernel. This result can be transfered to Block Graph Deletion as well. Next, we present a kernelization algorithm for the Strictly Chordal Editing problem, producing a reduced instance with $O(k^{4})$ vertices. Eventually, we discuss how this approach leads to $O(k^{3})$ vertex-kernels for both Strictly Chordal Completion and Strictly Chordal Deletion. Above all, our purpose is to exhibit general techniques that might, we hope, be extended to kernelizations for edge modification problems towards other graph classes. Several such algorithms, e.g., [4, 21] share the following feature. Very informally, the target class $\mathcal{G}$ admits a tree-like decomposition, in the sense that the vertices of any graph $H\in{\mathcal{G}}$ can be partitioned into clique modules (having the same neighborhood in $H$), and these modules can be mapped onto the nodes of a decomposition tree, the structure of the tree describing the adjacencies between modules. Therefore, if an arbitrary graph $G$ can be transformed into graph $H$ by at most $k$ edge additions or deletions, at most $2k$ modules can be affected by the modifications. By removing the affected nodes from the decomposition tree, we are left with several components (chunks) that correspond, in the initial graph $G$ as well as in $H$, to induced subgraphs that may be large but that already belong to the target class. Moreover, these chunks are attached to the rest of graph $G$ in a very regular way, through one or two nodes of the decomposition tree. The kernelization algorithms need to analyze these chunks and provide reduction rules, typically by ensuring a small number of nodes in the decomposition tree, plus the fact that each node corresponds to a module of small size.
Both the classes of block graphs and of strictly chordal graphs do not have exactly a tree-like decomposition, but still can be decomposed into structures than can be seen as a generalization of a tree. Our algorithms exploit these informal observations and provide the necessary reduction rules together with the combinatorial analysis for the kernel size.
Outline
The paper is organized as follows. We begin with some preliminary results and definitions, including the NP-completeness proofs of aforementioned problems (Section 2). We then describe a kernel with $O(k^{2})$ vertices for Block Graph Editing and Block Graph Deletion (Section 3), introducing the techniques that will be adapted to obtain an $O(k^{4})$ vertex-kernel for Strictly Chordal Editing (Section 4). We then explain how these techniques produce $O(k^{3})$ vertex-kernels for both Strictly Chordal Completion and Strictly Chordal Deletion (Section 5) and then conclude with some perspective and open problems (Section 6).
2 Preliminaries
We consider simple, undirected graphs $G=(V,E)$ where $V$ denotes the vertex set and $E\subseteq(V\times V)$ the edge set of $G$. We will sometimes use $V(G)$ and $E(G)$ to clarify the context. Given a vertex $u\in V$, the open neighborhood of $u$ is the set $N_{G}(u)=\{v\in V:\ uv\in E\}$. The closed neighborhood of $u$ is defined as $N_{G}[u]=N_{G}(u)\cup\{u\}$. Two vertices $u$ and $v$ are true twins if $N_{G}[u]=N_{G}[v]$. Given a subset of vertices $S\subseteq V$, $N_{G}[S]$ is the set $\cup_{v\in S}N_{G}[v]$ and $N_{G}(S)$ is the set $N_{G}[S]\setminus S$. We will omit the mention to $G$ whenever the context is clear. A subset of vertices $M\subseteq V$ is a module if for every vertices $x,y\in M$, $N(x)\setminus M=N(y)\setminus M$. The subgraph induced by $S$ is defined as $G[S]=(S,E_{S})$ where $E_{S}=\{uv\in E:\ u\in S,v\in S\}$. For the sake of readability, given a subset $S\subseteq V$ we define $G\setminus S$ as $G[V\setminus S]$. A subgraph $C$ is a connected component of $G$ if it is a maximal connected subgraph of $G$. A graph is biconnected if it is still connected after removing any vertex. A subgraph $C$ is a biconnected component of $G$ if it is a maximal biconnected subgraph of $G$. A set $S\subseteq V$ is a separator of $G$ if $G\setminus S$ is not connected. Given two vertices $u$ and $v$ of $G$, the separator $S$ is a $uv$-separator if $u$ and $v$ lie in distinct connected components of $G\setminus S$. Moreover, $S$ is a minimal $uv$-separator if no proper subset of $S$ is a $uv$-separator. Finally, a separator $S$ is minimal if there exists a pair $\{u,v\}$ such that $S$ is a minimal $uv$-separator.
Figure 1: The diamond, dart, gem and hole (cycle of length at least $4$)
Definition 1.
Given a graph $G=(V,E)$, a critical clique of $G$ is a set $K\subseteq V$ such that $G[K]$ is a clique, $K$ is a module and is inclusion-wise maximal under this property.
Notice that $K$ is a maximal set of true twins and that the set $\mathcal{K}(G)$ of critical cliques of any graph $G$ partitions its vertex set $V(G)$. This leads to the following definition.
Definition 2 (Critical clique graph).
Let $G=(V,E)$ be a graph. The critical clique graph of $G$ is the graph $\mathcal{C}(G)=(\mathcal{K}(G),E_{\mathcal{C}})$ with $E_{\mathcal{C}}=\{KK^{\prime}\;|\;\forall u\in K,\forall v\in K^{\prime},uv\in E\}$.
Block and strictly chordal graphs
Block graphs are graphs in which every biconnected component is a clique. They can also be characterized as chordal graphs that do not contain diamonds as induced subgraph [2] (see Figure 1). A natural generalization of block graphs are strictly chordal graphs, also known as block duplicate graphs, that are obtained from block graphs by adding true twins [28, 29, 23].
Theorem 1 ([33]).
Let $G=(V,E)$ be a graph. The following properties are equivalent:
  1. 1.
    $G$ is a strictly chordal graph,
  2. 2.
    The critical clique graph $\mathcal{C}(G)$ is a block graph,
  3. 3.
    $G$ does not contain any dart, gem or hole as an induced subgraph (see Figure 1),
  4. 4.
    $G$ is chordal and the minimal separators of $G$ are pairwise disjoint.
We consider the following problems:
Block Graph Editing Input: A graph $G=(V,E)$, an integer $k\in\mathbb{N}$ Question: Does there exist a set of pairs $F\subseteq(V\times V)$ of size at most $k$ such that the graph $H=(V,E\triangle F)$ is a block graph, with $E\triangle F=(E\setminus F)\cup(F\setminus E)$?
The Strictly Chordal Editing problem is defined similarly, requiring the graph $H=(V,E\triangle F)$ to be strictly chordal. The Block Graph Completion and Strictly Chordal Completion (resp. Block Graph Deletion and Strictly Chordal Deletion) problems are defined similarly by requiring $F$ to be disjoint from (resp. included in) edge set $E$. Given a graph $G=(V,E)$, a set $F\subseteq(V\times V)$ such that the graph $H=(V,E\triangle F)$ belongs to the target graph class is called an edition of $G$. When $F$ is disjoint from $E$ (resp. included in $E$) it is called a completion (resp. a deletion) of $G$. For the sake of simplicity we use $G\triangle F$, $G+F$ and $G-F$ to denote the resulting graphs in all versions of the problems. In all cases, $F$ is optimal whenever it is minimum-sized. Given an instance $(G=(V,E),k)$ of any of the aforementioned problems, we say that $F$ is a $k$-edition (resp. $k$-completion, $k$-deletion) whenever $F$ is an edition (resp. completion, deletion) of size at most $k$. A vertex is affected by $F$ whenever it is contained in some pair of $F$. We say that an instance $(G=(V,E),k)$ is a yes-instance whenever it admits a $k$-edition (resp. $k$-completion, $k$-deletion). When applied to an instance $(I,k)$ of the problem, a reduction rule is said to be safe if $(I,k)$ is a yes-instance if and only if the reduced instance $(I^{\prime},k^{\prime})$ is a yes-instance.
We will use the following result that guarantees that any clique module of a given graph $G$ will remain a clique module in any optimal edition towards some hereditary class of graphs closed under true twin addition, in particular towards strictly chordal graphs.
Lemma 1 ([4]).
Let $\mathcal{G}$ be an hereditary class of graphs closed under true twin addition. For every graph $G=(V,E)$, there exists an optimal edition (resp. completion, deletion) $F$ into a graph of $\mathcal{G}$ such that for any two critical cliques $K$ and $K$’ either $(K\times K^{\prime})\subseteq F$ or $(K\times K^{\prime})\cap F=\emptyset$.
Notice in particular that Lemma 1 implies that whenever the target graph class is hereditary and closed under true twin addition one may reduce the size of critical cliques to $k+1$. We shall see Section 3 that this result does not apply to block graphs. However, we will still be able to bound the size of critical cliques by $O(k)$.
Join composition $\otimes$
We now turn our attention to the notion of join composition (operation $\otimes$) between two graphs. This operation will play an important part in the proofs of our reduction rules for all considered problems.
Definition 3.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two vertex-disjoint graphs and let $S_{1}\subseteq V_{1},S_{2}\subseteq V_{2}$. The join composition of $G_{1}$ and $G_{2}$ on $S_{1}$ and $S_{2}$, denoted $(G_{1},S_{1})\otimes(G_{2},S_{2})$, is the graph:
$(G_{1},S_{1})\otimes(G_{2},S_{2})=(V_{1}\cup V_{2},E_{1}\cup E_{2}\cup(S_{1}\times S_{2}))$
Lemma 2.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two disjoint chordal graphs, $S_{1}\subseteq V_{1}$ and $S_{2}\subseteq V_{2}$ be non-empty cliques of respectively $G_{1}$ and $G_{2}$ and $G=(G_{1},S_{1})\otimes(G_{2},S_{2})$. Then graph $G$ is chordal, and any minimal separator $S$ of $G$ is also a minimal separator of $G_{1}$ or $G_{2}$, except if $S=S_{i}$ for $i\in\{1,2\}$ and $S_{i}$ is strictly included in a maximal clique of $G_{i}$.
Proof.
The graph $G$ is chordal since any cycle of $G$ is either contained in one of $G_{1}$, $G_{2}$ or, if it intersects both of them, contains at least three vertices of $S_{1}\cup S_{2}$, which is a clique in $G$. In both cases, if the cycle has four or more vertices then it contains a chord.
Any minimal separator $S$ of chordal graph $G$ is the intersection of two maximal cliques $\Omega,\Omega^{\prime}$ of $G$, and moreover $S$ is a minimal $ab$-separator for any $a\in\Omega\setminus S$, $b\in\Omega^{\prime}\setminus S$ — in particular, $S$ is strictly included in each of the cliques (see, e.g., Theorem 7 and Lemma 5 in [22]). By construction of $G$, any of its maximal cliques is either a maximal clique of $G_{1}$ or of $G_{2}$, or it is exactly $S_{1}\cup S_{2}$.
If $S=\emptyset$, then graph $G$ is not connected. This only happens if one of the graphs $G_{i}$ is not connected, since neither $S_{1}$ nor $S_{2}$ are empty. In this case, $S$ is a minimal separator of $G_{1}$ or $G_{2}$.
From now on, we assume that $S$ is not empty. Consider first the case where none of $\Omega,\Omega^{\prime}$ equals $S_{1}\cup S_{2}$. Then both $\Omega,\Omega^{\prime}$ are contained in the same $G_{i}$, for $i\in\{1,2\}$ (otherwise their intersection $S$ would be empty). Assume w.l.o.g they are both in $G_{1}$, we claim that $S=\Omega\cap\Omega^{\prime}$ is a minimal separator of $G_{1}$. Take $a\in\Omega\setminus S$ and $b\in\Omega^{\prime}\setminus S$. Since $S$ separates $a$ and $b$ in $G$, it also separates them in $G_{1}$, which is a subgraph of $G$. But $S$ is in the common neighborhood of $a$ and $b$ in $G_{1}$, thus $S$ is minimal among the $ab$-separators of $G_{1}$, and the conclusion follows.
We are left with the case when $\Omega=S_{1}\cup S_{2}$ and $\Omega^{\prime}$ is contained in one of the two graphs $G_{i}$, say in $G_{1}$. Therefore $S\subseteq S_{1}$. If this inclusion is strict, we claim that $S$ is also a minimal separator of $G_{1}$. Indeed $S$ is a minimal separator in $G$ for any $a\in\Omega\setminus S$ and $b\in\Omega^{\prime}\setminus S$, therefore we can choose $a\in S_{1}\setminus S$. Again $S$ separates $a$ and $b$ in the subgraph $G_{1}$ of $G$, and since $S$ is in the common neighborhood of $a$ and $b$ in $G_{1}$, it is necessarily a minimal separator of $G$. We deduce that, if $S$ is not a minimal separator of $G_{1}$, we must have $S=S_{1}$. We conclude that $S$ is strictly contained in the maximal clique $\Omega^{\prime}$ of $G_{1}$, which proves our lemma. ∎
We now present sufficient conditions on the sets $S_{1}$ and $S_{2}$ so that if the graphs $G_{1}$ and $G_{2}$ are block or strictly chordal, then so is the graph $(G_{1},S_{1})\otimes(G_{2},S_{2})$.
Lemma 3.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two disjoint block graphs and let $S_{1}\subseteq V_{1},S_{2}\subseteq V_{2}$. The graph $G=(G_{1},S_{1})\otimes(G_{2},S_{2})$ is a block graph if for $i\in\{1,2\}$, $S_{i}$ is a single vertex or a maximal clique of $G_{i}$.
Proof.
Observe that minimal separators of block graphs are single vertices since the biconnected components of block graphs are cliques. From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ (resp. $S_{2}$) if it is a not maximal clique of $G_{1}$ (resp. $G_{2}$). Hence minimal separators of $G$ are single vertices and it follows that $G$ is a block graph. ∎
Lemma 4.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two disjoint strictly chordal graphs and let $S_{1}\subseteq V_{1},S_{2}\subseteq V_{2}$. The graph $G=(G_{1},S_{1})\otimes(G_{2},S_{2})$ is strictly chordal if for $i\in\{1,2\}$, $S_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$.
Proof.
From Lemma 2, $G$ is chordal and its minimal separators of are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ or $S_{2}$ if they are not maximal cliques. If $S_{i}$ intersects exactly one maximal clique, $i\in\{1,2\}$, it is clear that it does not intersect any minimal separator of $G_{i}$. If $S_{i}$ is a critical clique, we claim that either $S_{i}$ is a minimal separator of $G_{i}$ or does not intersect any minimal separator of $G_{i}$. Indeed, if there is a minimal (clique) separator $S$ of $G_{i}$ such that $S_{i}\subset S$, then since $S_{i}$ is a critical clique, there exist $u\in S_{i},v\in S\setminus S_{i}$ and $w\in V(G_{i})\setminus S$ such that $w$ is adjacent to exactly one of $u$ and $v$. Suppose w.l.o.g. that $w$ is adjacent to $u$ and not to $v$, then $u$ is in a minimal $vw$-separator that intersects $S$ and is not equal to it. This is a contradiction since $G_{i}$ is strictly chordal and from Theorem 1 its minimal separators are pairwise disjoint. It follows that the minimal separators of $G$ are pairwise disjoint and by Theorem 1 $G$ is strictly chordal. ∎
In the proofs of some rules for the kernel for Strictly Chordal Editing and its variants, we will have to do a join composition of some graphs $G_{2},\ldots,G_{r}$ on a graph $G_{1}$ with the same set of vertices $S_{1}\subseteq V(G_{1})$. The following observation guarantees us that this operation is possible.
Observation 1.
Let $G_{1}=(V_{1},E_{1}),\dots,G_{r}=(V_{r},E_{r})$ be disjoint strictly chordal graphs and for each $i\in\{1,\dots,r\}$ let $S_{i}\subseteq V_{i}$ be a critical clique, a maximal clique or a set that intersects exactly one maximal clique of $G_{i}$. If $S_{1}$ is a critical clique or intersects exactly one maximal clique of $G_{1}$, then $S_{1}$ is a critical clique or intersects exactly one maximal clique of $(G_{1},S_{1})\otimes(G_{2},S_{2})$. This implies by Lemma 4 that $(G_{1},S_{1})\otimes(\bigcup_{2\leqslant i\leqslant r}G_{i},\bigcup_{2\leqslant i\leqslant r}S_{i})$ is strictly chordal (see Figure 2).
Figure 2: Example of a graph $G=(G_{1},S_{1})\otimes(\bigcup_{2\leqslant i\leqslant 4}G_{i},\bigcup_{2\leqslant i\leqslant 4}S_{i})$ as constructed in Observation 1. There, $S_{1}$ intersects exactly one maximal clique of $G_{1}$ and is a critical clique of $G$; $S_{2}$ intersects exactly one maximal clique of $G_{2}$; $S_{3}$ is a maximal clique of $G_{3}$; $S_{4}$ is a critical clique of $G_{4}$.
Spanning block subgraph
The following result is crucial to bound the size of the kernels for all variants of the studied modification problems.
Lemma 5.
Let $G=(V,E)$ be a connected block graph, a set $A\subseteq V(G)$ and $T_{A}$ be a minimal connected induced subgraph of $G$ that spans all vertices of $A$. Denote by $f(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. The following holds:
  1. (i)
    The subgraph $T_{A}$ is unique,
  2. (ii)
    $|f(T_{A})\setminus A|\leqslant 3\cdot|A|$,
  3. (iii)
    The graph $T_{A}\setminus(A\cup f(T_{A}))$ contains at most $2\cdot|A|$ connected components.
Proof.
A convenient way to represent the tree-like structure of block graph $G$ is its block-cut tree $T_{BC}(G)$. Recall that the block-cut tree of a graph $G$ has two types of nodes: the block nodes correspond to blocks of $G$ (i.e, biconnected components which, in our case, are precisely the maximal cliques of $G$) and the cut nodes correspond to cut-vertices of $G$. We put an edge between a cut node and a block node in $T_{BC}(G)$ if the corresponding cut-vertex belongs to the corresponding block of $G$.
Each vertex of $G$ that is not a cut-vertex belongs to a unique block of $G$. Therefore we can map each vertex $v$ of $V(G)$ on a unique node $n(v)$ of $T_{BC}(G)$ as follows: if $v$ is a cut-vertex, we map it on its corresponding cut node in $T_{BC}(G)$, otherwise we map it on the block node corresponding to the unique block of $G$ containing $v$. Observe that, for any two vertices $u,v\in A$, any (elementary) path $P$ from $u$ to $v$ in $G$ corresponds to the path $P_{BC}$ from $n(u)$ to $n(v)$ in $T_{BC}(G)$. In particular, $P$ contains all vertices corresponding to cut nodes of $P_{BC}$, therefore $T_{A}$ must contain all these vertices. Altogether, the vertices of $T_{A}$ are precisely the vertices of $A$, plus all cut-vertices of $G$ corresponding to the cut nodes of such paths, implying the unicity of $T_{A}$.
Let $T_{BC}(A)$ denote the subtree of $T_{BC}(G)$ spanning all nodes of $n(A)=\{n(a)\mid a\in A\}$. We count the vertices of $f(T_{A})\setminus A$. If $|A|=2$ then $f(T_{A})\setminus A=\emptyset$. Suppose now that $|A|\geqslant 3$ and let $a^{\prime}\in f(T_{A})\setminus A$. By the previous observation, it is a cut-vertex of $G$, so $n(a^{\prime})$ is a cut node of $T_{BC}(A)$. Let $v$ be a neighbor of $a^{\prime}$ in $T_{A}$. By construction of $T_{A}$ and $T_{BC}(A)$, there is a block node $b$ adjacent to $n(a^{\prime})$ in $T_{BC}(A)$, such that $v$ and $a^{\prime}$ are in the maximal clique of $G$ corresponding to node $b$. Moreover, $v$ is in $A$, or $v$ is a cut-vertex of $G$ such that the cut node $n(v)$ is adjacent to $b$ in $T_{BC}(A)$. Hence we have:
  1. 1.
    $n(a^{\prime})$ is of degree at least 3 in $T_{BC}(A)$, or
  2. 2.
    $n(a^{\prime})$ is the neighbor of a block node $b$ of degree at least three in $T_{BC}(A)$, or, if none of these hold, then
  3. 3.
    $n(a^{\prime})$ has exactly two neighbors $b$ and $b^{\prime}$ in $T_{BC}(A)$, the corresponding maximal cliques of $G$ contain at least $p$ vertices of $A$, where $p$ plus the degrees of $b$ and $b^{\prime}$ in $T_{BC}(A)$ is at least three.
Let $k$ be the number of leaves of $T_{BC}(A)$. Observe that for any leaf $l$ of $T_{BC}(A)$, there is some $a\in A$ such that $n(a)=l$. Choose for each leaf $l$ a unique vertex $a_{l}\in A$ such that $n(a_{l})=l$, we call $a_{l}$ a leaf vertex. Note that the number of vertices $a^{\prime}\in f(T_{A})$ corresponding to the first two items of the above enumeration is upper bounded by $3k$. Indeed, $n(a^{\prime})$ is incident to an edge of $T_{BC}(A)$, having an end node of degree at least 3. One can easily check that, in any tree of $k$ leaves, the number of such edges is bounded by $3k$ (this can be shown by induction on the number of leaves of the tree, adding a new leaf node at a time). We count now the vertices $a^{\prime}\in f(T_{A})$ of the third type. By the third item, $a^{\prime}$ has at least one neighbor $a\in A$ in graph $T_{A}$, such that $a$ is not a leaf vertex. Observe that $a$ can be in the neighborhood of at most two vertices $a^{\prime}\in f(T_{A})$ of this third type. Altogether it follows that $|f(T_{A})\setminus A|\leqslant 3k+2(|A|-k)\leqslant 3\cdot|A|$.
To prove the third item, the number of components of $T_{A}\setminus(A\cup f(T_{A}))$, we visualize again the situation in $T_{BC}(A)$. Recall that any node of degree at least three in $T_{BC}(A)$ is either a cut node, in which case it is in $f(T_{A})$ (the first case of the enumeration above) or a block node, but then all its neighbors in $T_{BC}(A)$ correspond to vertices of $f(T_{A})$ (the second case of the enumeration above). Therefore the components of $T_{A}\setminus f(T_{A})$ correspond to the components of $T_{BC}(A)$ after removal of all nodes of degree at least $3$. Hence the number of such components is upper bounded by $2k$. Removing the leaf vertices of $A$ does not increase the number of components, and the removal of each other vertex of $A$ increases the number of components by at most one. Thus $T_{A}\setminus(A\cup f(T_{A}))$ has at most $2\cdot|A|$ components, concluding the proof of our lemma. ∎
2.1 Hardness results
We first show the NP-completeness of Block Graph Deletion and Block Graph Editing by giving a reduction from Cluster Deletion and Cluster Editing, known to be NP-complete [32, 37, 34]. Notice that Block Graph Completion admits a polynomial-time algorithm due to the very nature of block graphs: any biconnected component of the input graph (computable in linear time [26]) must be turned into a clique.
Lemma 6.
Block Graph Deletion and Block Graph Editing are NP-Complete.
Proof.
We first prove that Block Graph Deletion is NP-Complete by providing a reduction from Cluster Deletion. A graph is a cluster graph if it does not contain any induced path on three vertices (so-called $P_{3}$). Given an instance $(G=(V,E),k)$ of Cluster Deletion, we construct an instance of Block Graph Deletion by adding a universal vertex $u$ adjacent to all vertices of $V$. Let $(G^{\prime}=(V^{\prime},E^{\prime}),k)$ be the produced instance.
We show that the graph $G$ admits a $k$-deletion into a cluster graph if and only if $G^{\prime}$ admits a $k$-deletion into a block graph. Suppose first that there is a $k$-deletion $F$ of $G$ into a cluster graph. The graph $G-F$ is a graph without any $P_{3}$ as induced subgraph. Now consider the graph $H^{\prime}=G^{\prime}-F$. By construction $H^{\prime}[V]$ contains no $P_{3}$, so $H^{\prime}$ is chordal and contains no diamond, thus $H^{\prime}$ is a block graph.
Now suppose that there exists a $k$-deletion $F^{\prime}$ of $G^{\prime}$ into a block graph. We claim that either $F^{\prime}$ does not affect the universal vertex $u$ or we can construct a $k$-deletion $F^{*}$ from $F^{\prime}$ that does not affect $u$. To support this claim, assume that $F^{\prime}$ contains at least one edge incident to $u$.
First, suppose that $H^{\prime}=G^{\prime}-F^{\prime}$ is not connected. Let $C$ be the connected component that contains $u$. For any other connected component $C^{\prime}$ of $H^{\prime}$, take a vertex $v\in V(C^{\prime})$ and construct the deletion set $F^{\prime\prime}=F^{\prime}\setminus\{(u,v)\}$. By Lemma 3, $(C,\{u\})\otimes(C^{\prime},\{v\})$ is a block graph, and the other connected components of $H^{\prime}$ are unchanged. Hence $G^{\prime}-F^{\prime\prime}$ is a block graph and since $|F^{\prime\prime}|<|F^{\prime}|$, $F^{\prime\prime}$ is a $k$-deletion. We iterate this construction until we get an intermediate solution $F^{*}_{1}$ such that $H^{*}_{1}=G^{\prime}-F^{*}_{1}$ is connected.
If $F^{*}_{1}$ does not affect $u$ we set $F^{*}=F^{*}_{1}$ and we are done. Otherwise there is at least one vertex $v$ not adjacent to $u$ in $H^{*}_{1}$. Since $H^{*}_{1}$ is connected we can choose $v$ at distance exactly $2$ from $u$. Since $H^{*}_{1}$ is a block graph we have $N(u)\cap N(v)=\{w\}$. Consider the biconnected component $C$ containing $v$ and $w$, and let $C^{\prime}=C\setminus\{w\}$. Since $H^{*}_{1}$ is a block graph, $C$ (and thus $C^{\prime}$) is a clique. Let $K_{u}$ and $K_{v}$ be the connected components of $H^{*}_{1}-(\{w\}\times C^{\prime})$ that contain the vertices $u$ and $v$ respectively, they are block graphs by heredity. By Lemma 3 the graph $H^{*}_{2}=(K_{u},\{u\})\otimes(K_{v},\{C^{\prime}\})$ is a block graph. Observe that ($(\{u\}\times C^{\prime})\subseteq F^{*}_{1}$ since $u$ is an universal vertex and let $F^{*}_{2}$ be the deletion set such that $H^{*}_{2}=G^{\prime}-F^{*}_{2}$. By construction we can observe that $|F^{*}_{2}|=|F^{*}_{1}|$, hence $F^{*}_{2}$ is a $k$-deletion of $G^{\prime}$ into a block graph. We iterate this construction until we get a $k$-deletion $F^{*}$ of $G$ that does not affect $u$.
Finally, $H=G-F^{*}$ does not contains any $P_{3}$ or else there would be a diamond in $G^{\prime}-F^{*}$ with the universal vertex $u$, thus $H$ is a cluster graph.
The same reduction can be done for Block Graph Editing, one needs only to observe that there are no added edges incident to $u$, and we can use the same argument as above on the deleted edges: if there is no deleted edge incident to $u$ we are done, otherwise we construct another set of deleted edges of equal size, but containing fewer edges incident to $u$. If in the construction we have to delete an edge that was added by the edition, it is clear that removing this edge from the edition set result in a strictly smaller edition set. ∎
The NP-completeness of Strictly Chordal Completion follows directly from the proof of NP-completeness of 3-Leaf Power Completion from [16, Theorem 3].
Lemma 7.
Strictly Chordal Completion is NP-complete.
We show the NP-completeness of Strictly Chordal Editing and Strictly Chordal Deletion by giving a reduction from Cluster Editing and Cluster Deletion, respectively.
Lemma 8.
Strictly Chordal Editing and Strictly Chordal Deletion are NP-complete.
Proof.
Given an instance $(G=(V,E),k)$ of Cluster Editing, we construct an instance of Strictly Chordal Editing by adding a clique $U=\{u_{1},\dots,u_{k+1}\}$ of size $k+1$ adjacent to all vertices of $V$, and for each vertex $x$ in $V$, $k+1$ vertices $\{v^{x}_{1},\dots,v^{x}_{k+1}\}$ adjacent only to $x$. Let $(G^{\prime}=(V^{\prime},E^{\prime}),k)$ be the produced instance. We show that the graph $G$ admits a $k$-edition into a cluster graph if and only if $G^{\prime}$ admits a $k$-edition into a strictly chordal graph. Suppose first that there is a $k$-edition $F$ of $G$ into a cluster graph. The graph $G\triangle F$ is a graph without any $P_{3}$ as induced subgraph. Consider the graph $H^{\prime}=G^{\prime}\triangle F$. By construction $H^{\prime}[V]$ contains no $P_{3}$, so $H^{\prime}$ is chordal and contains neither gems nor darts since these obstructions contain an induced $P_{3}$ and a vertex adjacent to every vertex of this $P_{3}$. By Theorem 1, it follows that $H^{\prime}$ is strictly chordal. Now suppose that there exists a $k$-edition $F^{\prime}$ of $G^{\prime}$ into a strictly chordal graph $H^{\prime}=G^{\prime}\triangle F^{\prime}$. We claim that $H=G\triangle F^{\prime}$ is a cluster graph. By contradiction, suppose that $H=H^{\prime}[V]$ contains a $P_{3}$$\{x,y,z\}$ where $x,z$ are the ends of the path. Then, there exist $i,j\in\{1,\dots,k+1\}$ such that $\{x,y,z,u_{i},v^{y}_{j}\}$ forms a dart in $H^{\prime}$, contradicting that $H^{\prime}$ is strictly chordal, and thus $H$ is a cluster graph.
The same reduction can be done from Cluster Deletion to Strictly Chordal Deletion. This concludes the proof. ∎
3 Kernelization algorithm for Block Graph Editing
We begin this section by providing a high-level description of our kernelization algorithm. A similar technique will be applied for modification problems towards strictly chordal graphs but on the critical clique graph rather than the original one. Let us consider a positive instance $(G=(V,E),k)$ of Block Graph Editing, $F$ a suitable solution and $H=G\triangle F$. Since $|F|\leqslant k$, we know that at most $2k$ vertices of $H$ may be affected vertices. Let $A$ be the set of such vertices, $T$ the minimum induced subgraph of $H$ that spans all vertices of $A$ and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. From Lemma 5 we have $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$. We will define the notion of BG-branch, corresponding to subgraphs of $G$ that induce block graphs. We will focus our interest on two types of BG-branches: the ones that are connected to the rest of the graph by only one (cut) vertex, called $1$-BG-branches, and the ones that are connected to the rest of the graph by exactly two non adjacent vertices, called $2$-BG-branches. We will prove that we can keep only the cut-vertices of $1$-BG-branches and reduce the $2$-BG-branches containing more than $k+3$ vertices. We will see that there are two kinds of connected components remaining in the graph $H\setminus(A\cup A^{\prime})$, the ones adjacent to maximal cliques of vertices of $A\cup A^{\prime}$, corresponding to critical cliques (which can be bounded to a linear number of vertices) and the ones adjacent to two non adjacent vertices of $A\cup A^{\prime}$ (which correspond to $2$-BG-branches in $H$). Since $H$ is a block graph, we know that there is at most $|A\cup A^{\prime}|=O(k)$ maximal cliques of vertices of $A\cup A^{\prime}$. From Lemma 5 we have that there are at most $4k$ connected components in $T\setminus(A\cup A^{\prime})$, and each one corresponds to a $2$-BG-branch in the graph $H$. It remains that $H\setminus(A\cup A^{\prime})$ contains a linear number of connected components, each one containing $O(k)$ vertices. Altogether, the graph $H$ contains $O(k^{2})$ vertices.
3.1 Classical reduction rules
We first give classical reduction rules when dealing with modification problems. Notice that Lemma 1 cannot be directly applied to Block Graph Editing and Block Graph Deletion since block graphs are not closed under true twin addition. Hence one cannot directly reduce the size of critical cliques to $k+1$ as mentioned Section 2. To circumvent this issue, we provide a slightly weaker reduction rule.
Rule 3.1.
Let $C$ be a block connected component of $G$. Remove $V(C)$ from $G$.
Rule 3.2.
Let $K\subseteq V$ be a set of true twins of $G$ such that $|K|>2k+2$. Remove $|K|-(2k+2)$ arbitrary vertices from $K$ in $G$.
Lemma 9.
Rules 3.1 and 3.2 are safe and can be computed in linear time.
Proof.
Rule 3.1 is safe since block graphs are hereditary and closed under disjoint union.
We now turn our attention to Rule 3.2. Let $G^{\prime}$ denote the graph obtained from the removal of $|K|-(2k+2)$ arbitrary vertices from $K$ in $G$ and $K^{\prime}=V(G)\setminus V(G^{\prime})$. Assume first that $G$ admits a $k$-edition $F$. Observe that $F$ is a $k$-edition of $G^{\prime}$ by heredity. Conversely, assume that $G^{\prime}$ admits a $k$-edition $F$ and let $H^{\prime}=G^{\prime}\triangle F$. Notice that since $F$ affects at most $2k$ vertices and since $|K\setminus K^{\prime}|=2k+2$, there exist two unaffected vertices $u$ and $v$ in $K\setminus K^{\prime}$. We claim that $H=G\triangle F^{\prime}$ is a block graph. Assume for a contradiction that $H$ contains an obstruction $W$. By construction, $W$ must intersect $K^{\prime}$ since otherwise $W$ would exist in $H^{\prime}$. Moreover, notice that $|W\cap K^{\prime}|\leqslant 2$, otherwise $W$ would contain a clique module with $3$ or $4$ vertices, which is impossible in a cycle or a diamond. Hence assume first that $|W\cap K^{\prime}|=\{a\}$: in this case, since $u$ is unaffected we have $N_{H}[a]\setminus K^{\prime}=N_{H^{\prime}}[u]$ and thus the set $(W\setminus\{a\})\cup\{u\}$ induces an obstruction in $H^{\prime}$, a contradiction. Similarly, if $|W\cap K^{\prime}|=\{a,b\}$ the set $(W\setminus\{a,b\})\cup\{u,v\}$ induces an obstruction in $H^{\prime}$, leading once again to a contradiction. Finally, since the biconnected components can be computed in $O(n+m)$ time [26], since block graphs can be recognized in $O(n+m)$ time and since true twins can be detected in $O(n+m)$ time (using modular decomposition [36]), both rules can be computed in linear time. ∎
3.2 Reducing Block Graph branches
Let $G=(V,E)$ be a graph, a BG-branch is a connected induced subgraph $B$ of $G$ such that $B$ is a block graph. A vertex $v\in V(B)$ is an attachment point of $B$ if $N_{G}(v)\setminus V(B)\neq\emptyset$. A BG-branch is a $p$-BG-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the subgraph of $B$ in which all attachment points have been removed. We first define a simple rule that allows to remove all the vertices of a $2$-BG-branch except for its attachment point.
Rule 3.3.
Let $B$ be a $1$-BG-branch of $G$ such that $B^{R}$ is connected and not empty. Remove $V(B^{R})$ from $G$.
Lemma 10.
Rule 3.3 is safe.
Proof.
Let $B$ be such a $1$-BG-branch and $G^{\prime}=G\setminus V(B^{R})$. Let $F$ be a $k$-edition of $G$ into a block graph. Since the class of block graphs is hereditary, $(G\triangle F)\setminus V(B^{R})$ is a block graph. It follows that $G^{\prime}\triangle F$ is a block graph and $F$ is a $k$-edition of $G^{\prime}$ into a block graph.
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$ into a block graph. Let $u$ be the attachment point of $B$ in $G$. Observe that since $B^{R}$ is connected and not empty, $u$ is not a cut-vertex of the block graph $B$ and its neighborhood in $B$ is either a maximal clique or a single vertex. Since $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $B$ are block graphs and $N_{B}(u)$ is a single vertex or a maximal clique of $B$, Lemma 3 ensures that $H=(H^{\prime},\{u\})\otimes(B^{R},N_{B}(u))$ is a block graph. We can observe that $H=G\triangle F^{\prime}$, hence $F^{\prime}$ is a $k$-edition of $G$ into a block graph. This concludes the proof. ∎
We now deal with the $2$-BG-branches, which require more technical skills. A $2$-BG-branch $B$ is clean if $B^{R}$ is connected and if the two attachment points are not adjacent. For a clean $2$-BG-branch $B$, a min-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that the attachment points $p_{1}$ and $p_{2}$ are not in the same connected component of $B-M$ and $mc(B)$ is the size of a min-cut of $B$. A $2$-BG-branch $B$ is clean if $B^{R}$ is connected and if the two attachment points are not adjacent. For a $2$-BG-branch proper $B$, a minimum edge-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that the attachment points $p_{1}$ and $p_{2}$ are not in the same connected component of $B-M$. We denote by $mc(B)$ the size of a minimum edge-cut of $B$.
Lemma 11.
Let $(G,k)$ be a yes-instance of Block Graph Editing reduced by Rule 3.3, and $B$ a clean $2$-BG-branch of $G$. If $|V(B)|\geqslant k+3$ then there exists a $k$-edition $F$ of $G$ into a block graph such that the set of pairs of $F$ containing a vertex of $V(B^{R})$ is either empty or exactly a min-cut of $B$.
Proof.
Let $B$ be a clean $2$-BG-branch in $G$ with attachment points $p_{1},p_{2}$ such that $|V(B)|\geqslant k+3$. We can observe that $N_{B}(p_{1})$ and $N_{B}(p_{2})$ are maximal cliques or single vertices since $B^{R}$ is connected. Since $B$ is reduced by Rule 3.3, we can observe that the cut-vertices of $B$ are contained in any path $\pi$ from $p_{1}$ to $p_{2}$. Indeed, if there is a cut-vertex $x$ that is not in $\pi$, then $p_{1}$ and $p_{2}$ are in the same connected component of $B\setminus\{x\}$, implying that there is a connected component $K$ of $B\setminus\{x\}$ containing neither $p_{1}$ nor $p_{2}$. Since $B$ is a $2$-BG-branch, $N_{G}(K)=\{x\}$, hence $G[V(K)\cup\{x\}]$ induces a $1$-BG-branch of $G$ with attachment point $x$, a contradiction. Notice in particular that there is a unique shortest path $\pi_{s}$ between $p_{1}$ and $p_{2}$ that contain all cut-vertices of $B$. Moreover, since the only cut-vertices of $B$ are the ones in $\pi_{s}$, a vertex of $B$ that is not in $\pi_{s}$ lies in exactly one maximal clique of $B$ containing two consecutive vertices of $\pi_{S}$.
Figure 3: Illustration of the case where $p_{1}$ and $p_{2}$ are not in the same connected component of $H^{\prime}$ in the Lemma 11. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$.
Let $F$ be a $k$-edition of $G$ into a block graph, $H=G\triangle F$ and $H^{\prime}=H\setminus V(B^{R})$. First consider the case in which $p_{1}$ and $p_{2}$ are in two different connected components in $H^{\prime}$. Let $H^{\prime}_{1},H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $p_{1}$ and $p_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. It follows from Lemma 3 that the graph $H^{*}$ resulting from the disjoint union of
$(((B^{R},N_{B}(p_{1}))\otimes(H^{\prime}_{1},\{p_{1}\})),N_{B}(p_{2}))\otimes(H^{\prime}_{2},\{p_{2}\})$
and $H^{\prime}_{3}$ is a block graph (see Figure 3). Let $F^{*}$ be the edition such that $H^{*}=G\triangle F^{*}$. By construction, $F^{*}$ is a subset of $F$ with no pair containing a vertex of $V(B^{R})$.
Assume now that $p_{1}$ and $p_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between them. Note that its length is potentially $1$ if $p_{1}$ and $p_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{p_{1}=u_{1},u_{2},\dots,u_{r}=p_{2}\}$ (not necessarily induced) from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Since $B$ is a clean $2$-BG-branch of $G$ reduced by Rule 3.3, we know by the previous arguments that every cut-vertex of $B$ is in $\pi_{B}$. Note that all vertices in $\pi_{B}$ belong to the same biconnected component of $H$, which is a clique. Hence we can consider w.l.o.g. that $\pi_{B}$ is the path containing exactly $p_{1},p_{2}$ and the cut-vertices of $B$, which is the shortest path in $B$ between $p_{1}$ and $p_{2}$. Since the vertices of $\pi_{B}$ induce a path in $B$ (on at least $3$ vertices since $p_{1}$ and $p_{2}$ are not adjacent) and a clique in $B\triangle F$, then for every $i\in\{1,\dots,r-2\}$, we have $(u_{i},u_{r})\in F$. Any vertex $v\in B\setminus V(\pi_{B})$ is adjacent in $B$ to two vertices $w_{1},w_{2}\in V(\pi_{B})$ and if $v$ remains in the same biconnected component as $w_{1}$ and $w_{2}$ in $H$, then $(v,p_{1})\in F$ or $(v,p_{2})\in F$. Otherwise $F$ must contain $(v,w_{1})$ or $(v,w_{2})$. In any case, $F$ contains at least one pair for each vertex $v\in B\setminus\{u_{r-1},u_{r}\}$. It follows from the previous arguments that $|F|\geqslant|B|-2$. However, since $|B|\geqslant k+3$ we have $|F|\geqslant k+1$, contradicting that $F$ is a solution, therefore there is no such path $\pi_{B}$.
We can assume that if $p_{1}$ and $p_{2}$ are in the same component of $H^{\prime}$, $F$ contains an edge-cut of $B$ or there would be a path $\pi_{B}$ from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Let $M$ be a min-cut of $B$ and consider $B_{1}$ (resp. $B_{2}$) the connected component containing $p_{1}$ (resp. $p_{2}$) in $B\triangle M$. Both $B^{R}_{1}=B_{1}\setminus\{p_{1}\}$ and $B^{R}_{2}=B_{2}\setminus\{p_{2}\}$ are induced subgraphs (one of them possibly empty) of $B$ and block graphs by heredity. By Lemma 3, the following graph is a block graph:
$H^{*}=((H^{\prime},\{p_{1}\})\otimes(B^{R}_{1},N_{B_{1}}(p_{1})),\{p_{2}\})\otimes(B^{R}_{2},N_{B_{2}}(p_{2}))$
Figure 4: Illustration of the case where $p_{1}$ and $p_{2}$ are in the same connected component of $H^{\prime}$ and $F$ contains a min-cut of $B$ in Lemma 21. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$. The red dotted edges between $B_{1}$ and $B_{2}$ are edges of an edge-cut of $B$ deleted by editing.
Let $F^{*}$ the edition such that $H^{*}=G\triangle F^{*}$, we have $F^{*}=(F\setminus(V(B^{R})\times V(G)))\cup M$, since $M$ is a min-cut of $B$, $|F^{*}|\leqslant|F|$ and $F^{*}$ is a $k$-edition of $G$.
In all cases, there exists a $k$-edition $F^{*}$ of $G$ into a block graph such that the set pairs of $F^{*}$ containing a vertex of $V(B^{R})$ is either empty or exactly a min-cut of $B$. This concludes the proof. ∎
Rule 3.4.
Let $(G=(V,E),k)$ be an instance of Block Graph Editing and $B$ be a clean $2$-BG-branch of $G$ containing at least $k+3$ vertices with attachment points $p_{1},p_{2}$. Remove $V(B^{R})$ from $G$ and add a vertex $x$ adjacent to $p_{1}$ and $p_{2}$, a clique of size $min\{mc(B)-1,k\}$ adjacent to $p_{1}$ and $x$ and a clique of size $k$ adjacent to $p_{2}$ and $x$.
Lemma 12.
Rule 4 is safe.
Proof.
Let $G^{\prime}$ be the graph obtained from an application of Rule 3.4 on the $2$-BG-branch $B$ of graph $G$ with attachment points $p_{1},p_{2}$ and $S$ the vertices introduced by this rule. Observe that $B^{\prime}=G^{\prime}[S\cup\{p_{1}\}\cup\{p_{2}\}]$ is a clean $2$-BG-branch of $G^{\prime}$, $|V(B^{\prime})|\geqslant k+3$ and $mc(B^{\prime})=min\{mc(B),k+1\}$.
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 11 and $H=G\triangle F$. If $F$ contains a min-cut of $B$, observe that $mc(B^{\prime})=mc(B)\leqslant k$ and consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{m}$ containing $p_{i}$, where $F_{m}$ is a min-cut of $B^{\prime}$. By Lemma 3 the graph
$H^{\prime}=((H\setminus V(B^{R}),p_{1})\otimes(B_{1}^{\prime},N_{B^{\prime}}(p_{1})),p_{2})\otimes(B_{2}^{\prime},N_{B^{\prime}}(p_{2}))$
is a block graph. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $mc(B^{\prime})=mc(B)\leqslant k$, by construction $|F^{\prime}|=|F|$ , therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$. If $F$ does not contain a min-cut of $B$, then no vertices of $B^{R}$ is affected and $p_{1},p_{2}$ are not in the same connected component of $H\setminus V(B^{R})$. Let $H_{1}$ and $H_{2}$ be the connected components of $H\setminus V(B^{R})$ that contain respectively $p_{1}$ and $p_{2}$ and $H_{3}$ the remaining components of $H\setminus V(B^{R})$. By Lemma 3 the graph $H^{\prime}$ corresponding to the disjoint union of
$(((B^{\prime R},N_{B^{\prime}}(p_{1}))\otimes(H_{1},\{p_{1}\})),N_{B^{\prime}}(p_{2}))\otimes(H_{2},\{p_{2}\})$
and $H_{3}$ is a block graph. We can observe that $H^{\prime}=G^{\prime}\triangle F$, hence $F$ is a $k$-edition of $G^{\prime}$ into a block graph. The other way is symmetrical. ∎
3.3 Bounding the size of the kernel
To conclude our kernelization algorithm we first state that reduction rules involving BG-branches can be applied in polynomial time.
Lemma 13.
Rules 3.3 and 3.4 can be applied in polynomial time.
Proof.
We show that we can enumerate all $1$-BG-branches and $2$-BG-branches in polynomial time. Recall that recognition of block graphs can be computed in linear time. All $1$-BG-branches can be enumerated by removing a vertex $v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $v$. We proceed similarly to detect clean $2$-BG-branches by removing a pair of non-adjacent vertices $u,v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $u$ and $v$. ∎
Theorem 2.
Block Graph Editing admits a kernel with $O(k^{2})$ vertices.
Proof.
Let $(G=(V,E),k)$ be a reduced yes-instance of Block Graph Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. Let $A$ be the set of affected vertices of $H$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T$ be the connected minimal induced subgraph of $H$ that spans all affected vertices of $A$, and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. By Lemma 5, $|A^{\prime}\setminus A|\leqslant 3\cdot|A|\leqslant 6k$.
First, we can observe that if there is a cut-vertex $x\in V(H)$ that is not in $V(T)$, since $T$ is the connected minimal induced subgraph of $H$ that spans all affected vertices, there is a connected component $K$ of $H\setminus\{x\}$ that does not contain any affected vertex. Since $B=H[V(K)\cup\{x\}]$ is a block graph by heredity, it follows that $B$ induces a $1$-BG-branch with attachment point $x$ and can be reduced by Rule 3.3, a contradiction. It remains that every cut-vertex of $H$ lies in $V(T)$ and that every vertex of $V(H)\setminus V(T)$ is adjacent to a maximal clique of vertices or a single vertex of $V(T)$. Since the graph is reduced under Rule 3.3, the second case can not happen or else there would a $1$-BG-branch. Hence, there are two kinds of connected components in $H\setminus(A\cup A^{\prime})$, namely the ones that are adjacent to a maximal clique of vertices of $A\cup A^{\prime}$ and the ones that are adjacent to two non-adjacent vertices of $A\cup A^{\prime}$ (i.e., the ones that contain connected components of $T\setminus(A\cup A^{\prime})$). We say that these connected components are of type 1 or type 2 respectively (see Figure 5).
The type 1 connected components are cliques since $H$ is a block graph and have the same neighborhood in $H$, thus they are critical cliques reduced by Rule 3.2 and contains at most $22k+$ 2 vertices. Since $T$ is a block graph, there is at most $|A\cup A^{\prime}|\leqslant 8k$ maximal cliques of vertices of $A\cup A^{\prime}$, hence there are at most $8k\cdot(2k+2)$ vertices in connected components of type 1.
Now consider the connected components of type 2. We can observe that such a component $K$ together with its adjacent vertices $x,y\in A\cup A^{\prime}$ form a clean 2-$BG$-branch, and thus contains at most $32k+$ 3 vertices. By Lemma 5, $T\setminus(A\cup A^{\prime})$ contains at most $2\cdot|A|=4k$ connected components which is exactly the number of type 2 components, hence there are at most $4k\cdot(2k+1)$ vertices in connected components of type 2.
We conclude that $|V(G)|=O(k^{2})$. Finally we claim that the reduced instance can be computed in polynomial time. Indeed, all reduction rules can be exhaustively applied in polynomial time (Lemmas 9 and 13).
Figure 5: Illustration of a reduced instance in the proof of Theorem 2. Square nodes correspond to vertices of $T$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $A^{\prime}\setminus A$. Circle nodes correspond to critical cliques of $G$.
In this section we presented a kernel for Block Graph Editing. For the case of Block Graph Deletion, all the lemmata above are still true. In particular Lemma 11 holds by simply replacing ”$k$-edition” with ”$k$-deletion”, and the proof for the deletion variant is contained in the edition one. Using the same set of rules, we have a kernel with $O(k^{2})$ vertices for Block Graph Deletion.
Theorem 3.
Block Graph Deletion admits a kernel with $O(k^{2})$ vertices.
4 Kernelization algorithm for Strictly Chordal Editing
As in the previous section, we begin this section by providing a high-level description of our kernelization algorithm. We use the critical clique graph of strictly chordal graphs, which is a block graph (Theorem 1), to bound the number of vertices of a reduced instance. While the technique used to bound the size of the kernels for block graphs and strictly chordal graphs is similar, the structure and rules for the latter are more involved. We thus provide a stand-alone proof of this result for the sake of completeness.
Let us consider a positive instance $(G=(V,E),k)$ of Strictly Chordal Editing, $F$ a suitable solution and $H=G\triangle F$. Denote by $\mathcal{C}(H)$ the critical clique graph of $H$ as described in Definition Definition 2. Since $|F|\leqslant k$, we know that at most $2k$ critical cliques of $\mathcal{C}(H)$ may contain affected vertices. Let $A$ be the set of such critical cliques, $T$ the minimum subgraph of $\mathcal{C}(H)$ that spans all critical cliques of $A$ and $A^{\prime}$ the set of critical cliques of degree at least $3$ in $T$. Recall that $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$ (Lemma 5). We will define the notion of SC-branch, corresponding to well-structured strictly chordal subgraphs of $G$. We will focus our interest on two types of SC-branches: the ones that are connected to the rest of the graph by only one critical clique, called $1$-SC-branches, and the ones that are connected to the rest of the graph by exactly two critical cliques, called $2$-SC-branches. This time the $1$-SC-branches will only be reduced to two critical cliques. The connected components of the graph $T\setminus(A\cup A^{\prime})$ correspond to subgraphs of $2$-SC-branches and the length of such $2$-SC-branches will be bounded by $k+3$. We know that there are at most $4k$ such connected components, thus there is $O(k^{2})$ critical cliques in $T$. Finally, the connected components of the graph $\mathcal{C}(H)\setminus V(T)$ correspond to $1$-SC-branches or sets of connected $1$-SC-branches. The latter case will require a specific reduction rule to be dealt with. This implies that every critical clique or maximal clique of $T$ can have a linear number of $1$-SC-branches of $\mathcal{C}(H)\setminus V(T)$ adjacent to it. Altogether, the graph $\mathcal{C}(H)$ contains $O(k^{3})$ critical cliques, and each critical clique is of size at most $k+1$, hence the graph $G$ contains $O(k^{4})$ vertices.
From now on, we always assume that the considered optimal editions (resp. completions, deletions) satisfy Lemma 1.
4.1 Classical reduction rules
First, we give classical reduction rules when dealing with modification problems. The first rule is safe for any target graph class hereditary and closed under disjoint union. The second one comes from the fact that strictly chordal graphs are hereditary and closed under true twin addition, combined with Lemma 1.
Rule 4.1.
Let $C$ be a strictly chordal connected component of $G$. Remove $V(C)$ from $G$.
Rule 4.2.
Let $K\subseteq V$ be a set of true twins of $G$ such that $|K|>k+1$. Remove $|K|-(k+1)$ arbitrary vertices in $K$ from $G$.
Lemma 14 (Folklore, [4]).
Rules 4.1 and 4.2 are safe and can be applied in polynomial time.
4.2 Reducing Strictly Chordal branches
We now consider the main structure of our kernelization algorithm, namely SC-branches.
Definition 4 (SC-branch).
Let $G=(V,E)$ be a graph. We say that a subgraph $B$ of $G$ is a SC-branch if $V(B)$ is an union of critical cliques $K_{1},\dots,K_{r}\in\mathcal{K}(G)$ such that the subgraph of $\mathcal{C}(G)$ induced by $K_{1},\dots,K_{r}$ is a connected block graph.
We emphasize that our definition of a SC-branch $B$ is stronger than simply requiring $B$ to be an induced strictly chordal graph. For example, if $G$ is the dart graph, the subgraph obtained by removing the pendant vertex is strictly chordal, but it is not a SC-branch because the corresponding critical cliques do not form a block graph in $\mathcal{C}(G)$. Let $B$ be a SC-branch of graph $G$ and let $K_{1},\dots,K_{r}$ be the critical cliques of $G$ contained in $V(B)$. We say that $K_{i}$ is an attachment point of $B$ if $N_{G}(K_{i})\setminus V(B)\neq\emptyset$. A SC-branch $B$ is a $p$-SC-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the subgraph of $B$ in which all attachment points have been removed.
We first give structural lemmata on SC-branches that will be helpful to prove the safeness of our rules.
Lemma 15.
[…]
Proof.
First, we show that all sets $Q_{i}$ are cliques. Suppose that $Q_{i}$ is not a clique for some $1\leqslant i\leqslant r$. Let $x$ and $y$ be non adjacent vertices of $Q_{i}$ and $z\in P$. Since $G_{i}$ is connected, take a shortest path $\pi$ between $x$ and $y$ in $G_{i}$. The subgraph induced by the vertices $\{x,y,z\}$ and those of $\pi$ contains either a cycle of length at least $4$ if $z$ is not adjacent to any inner vertex of $\pi$, which is a contradiction since $B$ is a block graph, or a diamond with $z$ being one of its vertices of degree $3$. In the latter case, since $z$ is not in the same critical clique of $\mathcal{C}(G)$ as its true twin in the diamond, the critical cliques of $\mathcal{C}(G)$ that contain some vertices of this diamond also form a diamond in $\mathcal{C}(G)$. Such a diamond is formed by critical cliques of $G$ contained in $B$, contradicting the definition of a SC-branch. In all cases we have a contradiction, it remains that $Q_{i}$ is a clique in $G$.
Now suppose that $Q_{i}$, $1\leqslant i\leqslant r$ is neither a critical clique nor a maximal clique of $G_{i}$ and intersects at least two maximal cliques of $G_{i}$. Suppose first that $Q_{i}$ is not a module, then there exist $u,v\in Q_{i}$ and $x\in V(G_{i})\setminus Q_{i}$ adjacent to only $u$ or $v$. Since $Q_{i}$ is not a maximal clique of $G_{i}$, there exists $y\in V(G_{i})\setminus Q_{i}$ adjacent to $u$ and $v$. Together with $w\in P$, the vertices $\{u,v,w,x,y\}$ induce in $B$ a gem if $x$ and $y$ are adjacent and a dart otherwise, leading to a contradiction. If $Q_{i}$ is a module, since it is not a critical clique (i.e. a maximal clique module) there exist $u\in Q_{i}$ and $v\in V(G_{i})\setminus Q_{i}$ that lie in the same critical clique of $G_{i}$. Since $Q_{i}$ intersects two maximal cliques of $G_{i}$, there exist two non adjacent vertices $x,y\in V(G_{i})\setminus Q_{i}$ that are adjacent to both $u$ and $v$. Together with $w\in P$, the vertices $\{u,v,w,x,y\}$ induce a dart in $B$, leading to a contradiction. It remains that $Q_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$. ∎
Lemma 16.
Let $G=(V,E)$ be a graph and $B$ a SC-branch of $G$. Let $F$ be an optimal edition of $G$ that respects Lemma 1, and $H=G\triangle F$. For any attachment point $P$ of $B$, let $C$ be the critical clique of $H$ which contains $P$ and $C^{\prime}=C\setminus V(B^{R})$. If $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$, then $C^{\prime}$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}=H\setminus V(B^{R})$.
Proof.
Suppose that $C^{\prime}$ is not a critical clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Since $C^{\prime}\subseteq C$ is a module but not a critical clique, it is not a maximal clique module and thus there exist $u\in C^{\prime}$ and $v\notin C^{\prime}$ that lie in the same critical clique in $H^{\prime}$. Notice that there must exist a vertex $w\in N_{H}(C^{\prime})\cap V(B^{R})$ adjacent to exactly one of $u$ and $v$ in $H$, since otherwise we would have $N_{H}(u)\cap V(B)=N_{H}(v)\cap V(B)$ and $N_{H^{\prime}}(u)=N_{H^{\prime}}(v)$, meaning that $C$ is not a critical clique of $H$, a contradiction. Since $C^{\prime}$ intersects two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ if $w$ is not adjacent to neither $x$ nor $y$, a gem if it is adjacent to one of them, and if $w$ is adjacent to both $x$ and $y$, $\{w,x,v,y\}$ induce a $C_{4}$ in $H$. It remains that $C^{\prime}$ is a critical clique of $H^{\prime}$ or $C^{\prime}$ is included in exactly one maximal clique of $H^{\prime}$ if $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$. ∎
In the next result we show that there always exists an optimal edition that affects only the attachment point of a 1-SC-branch and its neighborhood. This will allow us to define our first reduction rule on SC-branches.
Lemma 17.
Let $G=(V,E)$ be a graph and $B$ a $1$-SC-branch of $G$ with attachment point $P$. There exists an optimal edition $F$ of $G$ such that:
  • The set of vertices of $B$ affected by $F$ is included in $P\cup N_{B}(P)$.
  • In $H=G\triangle F$ the vertices of $N_{B}(P)$ are all adjacent to the same vertices of $V(G)\setminus V(B^{R})$.
Proof.
[…]
$H^{*}=(\bigcup_{1\leqslant i\leqslant r}G_{i},\bigcup_{1\leqslant i\leqslant r}Q_{i})\otimes(H^{\prime},C^{\prime})$
If $P\cap N=\emptyset$, let $H^{*}$ be the strictly chordal graph corresponding to the disjoint union of $H^{\prime}$ and the graphs $G_{1},\dots,G_{r}$. Let $F^{*}$ be such that $H^{*}=G\triangle F^{*}$, in both cases by construction $|F^{*}|\leqslant|F|$ and the desired properties are verified. ∎
Figure 6: Illustration of the proof of Lemma 17 for $r=3$. The graph $H^{*}=(G_{1}\cup G_{2}\cup G_{3},Q_{1}\cup Q_{2}\cup Q_{3})\otimes(H^{\prime},C^{\prime})$ is strictly chordal by Lemma 4.
Rule 4.3.
Let $(G=(V,E),k)$ be an instance of Strictly Chordal Editing. If $G$ contains a $1$-SC-branch $B$ with attachment point $P$, then remove from $G$ the vertices of $V(B^{R})$ and add a clique $K$ of size $min\{|N_{B}(P)|,k+1\}$ adjacent to $P$.
Lemma 18.
Rule 4.3 is safe.
Proof.
Let $G^{\prime}$ be the graph obtained from an application of Rule 4.3 on $1$-SC-branch $B$ of graph $G$ with attachment point $P$. We denote by $K$ be the clique that replaced $B$. Let $F$ be a $k$-edition of $G$ and $H=G\triangle F$. By Lemma 17 we can assume that $F$ only affects vertices of $P\cup N_{B}(P)$ and all vertices of $N_{B}(P)$ are adjacent to the same vertices of $V(G)\setminus V(B^{R})$. This implies that if $|N_{B}(P)|>k$ no $k$-edition can affect these vertices. Let $C$ be the critical clique of $H$ which contains $P$ and $C^{\prime}=C\setminus V(B^{R})$. By Lemma 16, $C^{\prime}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. If $N_{B}(P)$ is adjacent to $C^{\prime}$ in $H$, consider the graph $H^{\prime}=(H_{a},C^{\prime})\otimes(K,K)$. Since $K$ is clique, by Lemma 4, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H^{\prime}$ corresponding to the disjoint union of $H_{a}$ and $K$. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F^{\prime}$ is a $k$-edition of $G^{\prime}$.
[…]
If the clique $K$ is adjacent to $C^{\prime}$ in $H^{\prime}$, consider the graph $H=(H_{a}^{\prime},C^{\prime})\otimes(\bigcup_{1\leqslant i\leqslant r}G_{i},\bigcup_{1\leqslant i\leqslant r}Q_{i})$. By Lemma 4 and Observation 1, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H$ corresponding to the disjoint union of $H_{a}^{\prime}$ and $G_{i}$ for $1\leqslant i\leqslant r$. Let $F$ be the edition such that $H=G\triangle F$, since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F$ is a $k$-edition of $G$. ∎
We now show a result that will allow us to define a rule bounding the number of $1$-SC-branches sharing the same neighborhood. In particular, if enough $1$-SC-branches share a neighborhood $N$, they force any editing of size at most $k$ to transform $N$ into a clique.
Lemma 19.
Let $(G=(V,E),k)$ be a yes-instance of Strictly Chordal Editing reduced by Rule 4.2. Let $B_{1},\dots,B_{l}$ be disjoint $1$-SC-branches of $G$ with attachment points $P_{1},\dots,P_{l}$ which have the same neighborhood $N$ in $G\setminus\bigcup_{i=1}^{l}V(B_{i})$ and form a disjoint union of cliques $Q_{1},\dots Q_{r}$ in $G[P_{1}\cup\dots\cup P_{l}]$. If $\Sigma_{i=1}^{l}|P_{i}|>2k+1$ then for every $k$-edition $F$ of $G$:
  1. (i)
    If $r=1$, $N$ is a clique of $H=G\triangle F$ and $N$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $H^{\prime}=H\setminus\bigcup_{i=1}^{l}V(B_{i})$,
  2. (ii)
    If $r>1$ and $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}>k$ then $N$ is a critical clique of $H=G\triangle F$ and $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}=H\setminus\bigcup_{i=1}^{l}V(B_{i})$.
Proof.
We prove the result by considering each item in separate cases.
Case (i) $r=1$:
We first prove that $N$ has to be clique in $H$. Suppose that $N$ is not a clique in $H$, then there exist two vertices $x,y\in N$ such that $xy\notin E(H)$. The attachment points are critical cliques, so by Rule 4.2, $|P_{i}|\leqslant k+1,i\in\{1,\dots,l\}$. We can observe that there is at most one critical clique $P_{j}$ such that $N_{B_{j}}(P_{j})=\emptyset$ (recall that $r=1$ and hence all attachment points lie in the same clique in $G[P_{1}\cup\dots\cup P_{l}]$). It follows that $|Q_{1}\setminus P_{j}|>k$, thus there exists a vertex $u$ in some $P_{t},t\neq j$ such that in $H$, $u$ is adjacent to a vertex $z\in N_{B_{t}}(P_{t})$ and the vertices $x$ and $y$. Since $|Q_{1}\setminus P_{t}|>k$, there exists a vertex $v\in Q_{1}\setminus P_{t}$ such that there is no pair of $F$ containing both $v$ and a vertex of $\{x,y,z,u\}$. The set $\{x,y,u,v,z\}$ induces a dart if neither $x$ nor $y$ is adjacent to $z$, a gem if one of them is adjacent to $z$, or if both of them are adjacent to $z$, $\{x,y,z,v\}$ induces a $C_{4}$ in $H$, a contradiction. It follows that $N$ has to be a clique.
We now show that $N$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $H^{\prime}$. Suppose that $N$ is neither a critical clique nor a maximal clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Suppose first that $N$ is not a module, there exist $u,v\in N$, $x\in V(H^{\prime})\setminus N$ adjacent to only $u$ or $v$. Since $N$ is not a maximal clique of $H^{\prime}$, there exist $y\in V(H^{\prime})\setminus N$ adjacent to $u$ and $v$. Together with a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$, the vertices $\{u,v,w,x,y\}$ induce in $H$ a gem if $x$ and $y$ are adjacent and a dart else, leading to a contradiction. If $N$ is a module and is not a critical clique, $N$ is not a maximal clique module, thus there exist $u\in N$, a vertex $v$ in the same module as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. Since $N$ intersects two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ since $w$ is only adjacent to $u$. It remains that if $r=1$, $N$ is a critical clique, a maximal clique or is included in exactly one maximal clique of $H^{\prime}$.
Case (ii) $r>1$ and $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}>k$:
We claim that $N$ is a clique in $H$. Suppose again for a contradiction that this is not the case, then there exist two vertices $x,y\in N$ such that $xy\notin E(H)$. For any pair of vertices $u_{i}\in Q_{i},u_{j}\in Q_{j},i\neq j$, the set $\{x,u_{i},y,u_{j}\}$ induces a $C_{4}$ in $H$ if neither $u_{j}$ nor $u_{i}$ is affected by $F$. It follows that the vertices of $Q_{i}$ or $Q_{j}$ must be affected by $F$. This implies that at most one $Q_{i}$ is unaffected by $F$, and since $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}>k$ we have $|F|>k$, a contradiction. Hence $N$ has to be a clique in $H$.
Next we show that $N$ is a critical clique of $H$. Suppose that $N$ is not a module in $H$, then there exist $x,y\in N$ and $z\in V(G)\setminus N$ adjacent to only one of the vertices $x$ or $y$, say w.l.o.g. $x$. For any pair of vertices $u_{i}\in Q_{i},u_{j}\in Q_{j},i\neq j$, if neither $u_{i}$ nor $u_{i}$ is affected by $F$, then at least one of them is not adjacent to $z$ and the set $\{x,y,u_{i},u_{j},z\}$ induces a dart if neither $u_{i}$ nor $u_{j}$ is adjacent to $z$ or a gem if one of them is adjacent to $z$. This implies that at most one $Q_{i}$ is unaffected by $F$, as before we have $|F|>k$, a contradiction. Hence $N$ has to be a module in $H$. If $N$ is not a critical clique of $H$, then it is strictly contained in some critical clique $N^{\prime}$. A vertex $x\in N^{\prime}\setminus N$ must have the same closed neighborhood than $N$ in $H$. If $x$ is not in some $P_{i}$ or $N_{B_{i}}(P_{i})$, then $F$ must affect all vertices of $P_{1}\cup\dots\cup P_{l}$, implying $|F|>k$, a contradiction. If $x$ is in some $P_{i}$ or $N_{B_{i}}(P_{i})$, then $F$ must affect all vertices of every $Q_{j}$ such that $P_{i}\nsubseteq Q_{j}$. Since $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}>k$, this implies $|F|>k$, a contradiction. It remains that $N$ is a critical clique in $H$.
Finally, since $N$ is a module in $H^{\prime}$, if it is not a critical clique, there exist $u\in N$, a vertex $v$ in the same critical clique as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. If $N$ intersects at least two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ since $w$ is only adjacent to $u$. It remains that $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}$. ∎
Rule 4.4.
Let $(G=(V,E),k)$ be an instance of Strictly Chordal Editing. Let $B_{1},\dots,B_{l}$ be disjoint $1$-SC-branches of $G$ with attachment points $P_{1},\dots,P_{l}$ which have the same neighborhood $N$ in $G\setminus\bigcup_{i=1}^{l}V(B_{i})$ and form a disjoint union of cliques $Q_{1},\dots Q_{r}$ in $G[P_{1}\cup\dots\cup P_{l}]$. Assume that $\Sigma_{i=1}^{l}|P_{i}|>2k+1$, then :
  • If $r=1$, remove the vertices $\bigcup_{i=1}^{l}V(B_{i})$, add two adjacent cliques $K$ and $K^{\prime}$ of size $k+1$ with neighborhood $N$ and a vertex $u_{K}$ adjacent to every vertex of $K$.
  • If $r>1$ and $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}>k$, remove the vertices $\bigcup_{i=1}^{l}V(B_{i})$ and add two non-adjacent cliques of size $k+1$ with neighborhood $N$.
Lemma 20.
Rule 4.4 is safe.
Proof.
Let $G^{\prime}$ be the graph obtained after the application of Rule 4.4 on $B_{1},\dots,B_{l}$, and $S$ the new vertices introduced by this rule. Notice that the subgraph $G_{S}^{\prime}=G^{\prime}[S]$ is strictly chordal and form two $1$-SC-branches that satisfy the conditions of Lemma 19. First, let $F$ be a $k$-edition of $G$, $H=G\triangle F$ and $H_{a}=H\setminus\bigcup_{i=1}^{l}V(B_{i})$. By Lemma 19 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}$. Therefore, by Lemma 4 and Observation 1 the graph $H^{\prime}=(H_{a},N)\otimes(G_{S}^{\prime},K\cup K^{\prime})$ is strictly chordal and the edition $F^{\prime}$ such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$ is of size at most $|F|\leqslant k$.
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$, $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $H_{a}^{\prime}=H^{\prime}\setminus S$. Notice that the subgraph $G_{b}=G[\bigcup_{i=1}^{l}V(B_{i})]$ is strictly chordal. By Lemma 19 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}^{\prime}$. Let $Q_{1},\dots,Q_{r}$ the cliques of $G_{b}$ adjacent to $N$, using arguments similar to the ones of Lemma 15, it is easy to see that each $Q_{i}$ is critical clique, a maximal clique or intersects exactly one maximal clique of $G_{b}$. Therefore, by Lemma 4 and Observation 1 the graph $H=(H_{a}^{\prime},N)\otimes(G_{b},Q_{1}\cup\dots\cup Q_{r})$ is strictly chordal and the edition $F$ such that $H=G\triangle F$ is of size at most $|F^{\prime}|\leqslant k$. ∎
4.3 Reducing the 2-SC-branches
Let $B$ be a $2$-SC-branch of a graph $G$ reduced by Rule 4.3, with attachment points $P_{1}$ and $P_{2}$. We say that $B$ is clean if $B^{R}$ is connected, and that the length of a clean $2$-SC-branch is the length of a shortest path between the two attachment points in $\mathcal{C}(B)$. We say that the min-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that $P_{1}$ and $P_{2}$ are not in the same connected component of $B-M$ and $mc(B)$ is the size of a min-cut of $B$. We can observe that a min-cut of $B$ contains the edges between a pair of consecutive critical cliques of a shortest path between $P_{1}$ and $P_{2}$ and the edges between one of these critical cliques and the critical cliques they have in their common neighborhood. The following observation will be useful to reduce $2$-SC-branches.
Observation 2.
Let $F$ be an optimal edition of $G$ and $F_{1}\subseteq F$. If $F_{2}$ is an optimal edition of the graph $G\triangle F_{1}$ , then $F_{1}\cup F_{2}$ is an optimal edition of $G$.
Lemma 21.
Let $(G=(V,E),k)$ be a yes-instance of Strictly Chordal Editing and $B$ a clean $2$-SC-branch of length at least $k+3$ with attachment points $P_{1},P_{2}$. There exists an optimal edition $F$ of $G$ such that:
  • If $P_{1}$ and $P_{2}$ are in the same connected component of $(G\triangle F)\setminus V(B^{R})$, then $F$ contains an edge-cut of $B$. This edge-cut is $(P_{1}\times N_{B}(P_{1}))$ or $(P_{2}\times N_{B}(P_{2}))$ or a min-cut of $B$.
  • In each case, the other vertices of $B$ affected by $F$ are included in $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$.
  • In $G\triangle F$ the vertices of $N_{B}(P_{1})$ (resp. $N_{B}(P_{2})$) are all adjacent to the same vertices of $V(G)\setminus V(B^{R})$.
Proof.
Let $F$ be an optimal edition of $G$, $H=G\triangle F$. Let $C_{1}$ (resp. $C_{2}$) be the critical clique of $H$ which contains $P_{1}$ (resp. $P_{2}$). Let $H^{\prime}=H\setminus V(B^{R})$, $C_{1}^{\prime}=C_{1}\setminus V(B^{R})$ and $C_{2}^{\prime}=C_{2}\setminus V(B^{R})$.
We first consider the case where $P_{1}$ and $P_{2}$ are in different connected component of $H^{\prime}$. Let $H^{\prime}_{1}$ and $H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $P_{1}$ and $P_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. For $i\in\{1,2\}$, Lemma 15 implies that $N_{B}(P_{i})$ is a clique in $G$ since $B^{R}$ is connected, and $N_{B}(P_{i})$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $B^{R}$. By Lemma 16, $C_{i}^{\prime}$ is a critical clique or intersects exactly one maximal clique of the connected component $H^{\prime}_{i}$.
By Lemma 4, the graph $H^{*}$ corresponding to the disjoint union of
$((B^{R},N_{B}(P_{1}))\otimes(H^{\prime}_{1},C_{1}^{\prime})),N_{B}(P_{2}))\otimes(H^{\prime}_{2},C_{2}^{\prime})$
and $H^{\prime}_{3}$ is strictly chordal (see Figure 7) and the edition $F^{*}$ such that $H^{*}=G\triangle F^{*}$ verifies the desired properties. Since $|F^{*}|\leqslant|F|$ by construction, we conclude that $F^{*}$ is optimal.
Figure 7: Illustration of the case where $B\triangle F$ is connected in Lemma 21.
Assume now that $P_{1}$ and $P_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between $c_{1}\in P_{1}$ and $c_{2}\in P_{2}$. Note that its length is potentially $1$ if $c_{1}$ and $c_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{c_{1}=u_{1},u_{2},\dots,u_{r}=c_{2}\}$ (not necessarily induced) from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$. Observe that such this path $\pi_{B}$ is of length at least $k+3$ since $B$ is of length at least $k+3$. There is no edge between the two paths in $G$, so $H[V(\pi_{B})\cup V(\pi_{H})]$ admits at least one induced cycle of length at least 4, which is a contradiction. We conclude that if $P_{1}$ and $P_{2}$ are in the same connected component of $H\setminus V(B^{R})$, then $F$ contains an edge-cut of $B$. Else there would be a path $\pi_{B}$ from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$.
Hence we can assume that $F$ contains an edge-cut of $B$. First, if $F_{1}=(P_{1}\times N_{B}(P_{1}))\subseteq F$, consider the graph $G_{1}=G\triangle F_{1}$. Observe that $B_{1}=B\setminus P_{1}$ is a $1$-SC-branch of $G_{1}$ with attachment point $P_{2}$. By Lemma 17 there exists an optimal edition $F_{2}$ of $G_{1}$ where the vertices of $B_{1}$ affected by $F_{2}$ are in $P_{2}\cup N_{B_{1}}(P_{2})$. By Observation 2, $F_{1}\cup F_{2}$ is an optimal edition and it verifies the desired properties. The same arguments can be used if $(P_{2}\times N_{B}(P_{2}))\subseteq F$.
Now, if $F$ contains neither $(P_{1}\times N_{B}(P_{1}))$ nor $(P_{2}\times N_{B}(P_{2}))$, let $F_{1}=F\cap(V(B^{R})\times V(B^{R}))$. There are two components in $B\triangle F_{1}$, say $B_{1}$ the one containing $P_{1}$ and $B_{2}$ one containing $P_{2}$. These connected components are $1$-SC-branches of $G_{1}=G\triangle F_{1}$ with attachment points $P_{1}$ and $P_{2}$. As before Lemma 17 applies on the $1$-SC-branches $B_{1}$ and $B_{2}$, thus there is an optimal edition $F_{2}$ of $G_{1}$ where the only vertices of $B_{1}$ and $B_{2}$ affected are included in $P_{1}\cup N_{B}(P_{1})$ and $P_{2}\cup N_{B}(P_{2})$. If $F_{1}$ is a min-cut of $B$, the edition $F_{1}\cup F_{2}$ is an optimal edition of $G$ that verifies the desired properties, else, replacing $F_{1}$ by a min-cut of $B$ result in a smaller edition that verifies the properties. ∎
Rule 4.5.
Let $(G=(V,E),k)$ be an instance of Strictly Chordal Editing and $B$ a clean $2$-SC-branch of $G$ of length at least $k+3$ with attachment points $P_{1},P_{2}$. Remove the vertices $V(B^{R})$, and add the following path of $k+2$ cliques:
$K_{min\{|N_{B}(P_{1})|,k+1\}}-K_{k+1}-K_{1}-K_{|mc(B)|}-K_{k+1}^{1}-\dots-K_{k+1}^{k-3}-K_{min\{|N_{B}(P_{2})|,k+1\}}$
where $K_{n}$ is the clique of size $n$ and $K_{min\{|N_{B}(P_{1})|,k+1\}}$ (resp. $K_{min\{|N_{B}(P_{2})|,k+1\}}$) is adjacent to $P_{1}$ (resp. $P_{2}$).
Lemma 22.
Rule 4.5 is safe.
Proof.
Let $G^{\prime}$ the graph obtained from an application of Rule 4.5 on the $2$-SC-branch $B$ of graph $G$ with attachment points $P_{1},P_{2}$ and $S$ the vertices introduced by this rule. Observe that $B^{\prime}=G^{\prime}[S\cup P_{1}\cup P_{2}]$ is a clean $2$-SC-branch of $G^{\prime}$ of length $k+3$ and $mc(B^{\prime})=min\{k+1,mc(B)\}$.
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 21 and $H=G\triangle F$. Observe that if $mc(B)\geqslant k+1$, then $F$ does not affect vertices of $V(B)$ other than $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$ and if $|N_{B}(P_{i})|>k,i\in\{1,2\}$ $F$ does not affect the vertices of $N_{B}(P_{i})$. Let $C_{i}$ be the critical clique of $H$ which contains $P_{i}$ and $C_{i}^{\prime}=C_{i}\setminus B^{R}$. By Lemma 16, $C^{\prime}_{i}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. Observe moreover that $N_{B^{\prime}}(P_{i})$ is critical clique of $B^{\prime R}$.
We now construct a $k$-edition $F^{\prime}$ of $G^{\prime}$. We consider first the tree cases where $F$ contains an edge-cut of $B$. If w.l.o.g. $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\cap F=\emptyset$. Observe that $(C_{1}^{\prime}\times N_{B^{\prime}}(P_{1})\subseteq F$ since $C_{1}$ is a critical clique of $H$. By Lemma 4 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B^{\prime R},N_{B^{\prime}}(P_{1}))$ is strictly chordal. If $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\subseteq F$ then $H^{\prime}$ the graph resulting of the disjoint union of $H_{a}$ and $B^{\prime R}$ is strictly chordal.If $F$ contains a min-cut of $B$, then consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{1}^{\prime}$ containing $P_{i}$, where $F_{1}^{\prime}$ is a min-cut of $B^{\prime}$. By Lemma 4 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B_{1}^{\prime},N_{B^{\prime}}(P_{1})),C_{2}^{\prime})\otimes(B_{2}^{\prime},N_{B^{\prime}}(P_{2}))$ is strictly chordal.
Otherwise, $P_{1}$ and $P_{2}$ are not in the same connected component of $H_{a}$ by Lemma 21. Let $H_{1}$ and $H_{2}$ be these components and $H_{3}$ be the remaining of $H_{a}$. By Lemma 4, the graph $H^{\prime}$ corresponding to the disjoint union of $((B^{\prime R},N_{B^{\prime}}(P_{1}))\otimes(H_{1},C_{1}^{\prime}),N_{B^{\prime}}(P_{2}))\otimes(H_{2},C_{2}^{\prime})$ and $H_{3}$ is strictly chordal.
In any of these four cases, let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$, by construction $|F^{\prime}|\leqslant|F|$, therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$.
The other way is symmetrical. ∎
4.4 Bounding the size of the kernel
We first state that reduction rules involving SC-branches can be applied in polynomial time.
Lemma 23.
Given an instance $(G=(V,E),k)$ of Strictly Chordal Editing, Rules 4.3, 4.4 and 4.5 can be applied in polynomial time.
Proof.
We rely on the linear time computation of the critical clique graph $\mathcal{C}(G)$ of $G$ [36] and the linear time recognition of block graphs. We show that we can enumerate all $1$-SC-branches and $2$-SC-branches in polynomial time. Since an attachment point is by definition a critical clique of $G$, one can detect $1$-SC-branches by removing a critical clique of $\mathcal{C}(G)$ and looking among the remaining connected components those that induce a connected block graph together with $P$ (in $\mathcal{C}(G)$). Considering a maximal set of such components together with $P$ gives a $1$-SC-branch $B$. We proceed similarly to detect clean $2$-SC-branches by removing a pair of critical cliques $P_{1},P_{2}$ of $\mathcal{C}(G)$, and looking among the remaining connected components those that induce a connected block graph (in $\mathcal{C}(G)$) together with $P_{1}$ and $P_{2}$. Such components together with $P_{1}$ and $P_{2}$ gives a clean $2$-SC-branch $B$. Since there is $O(|V(G)|)$ critical cliques, the result follows. ∎
Lemma 24 ([36]).
Let $G=(V,E)$ be a graph and $H=G\triangle\{uv\}$ for $u,v\in V$, then $|\mathcal{K}(H)|\leqslant|\mathcal{K}(G)|+2$.
Theorem 4.
Strictly Chordal Editing admits a kernel with $O(k^{4})$ vertices.
Proof.
Let $(G=(V,E),k)$ be a reduced yes-instance of Strictly Chordal Editing, $F$ a $k$-édition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. The graph $H$ is strictly chordal, thus the graph $\mathcal{C}(H)$ is a block graph. We first show that $\mathcal{C}(H)$ has $O(k^{3})$ vertices. We say that a critical clique of $\mathcal{C}(H)$ is affected if it contains a vertex affected by $F$. Let $A$ be the set of affected critical cliques of $\mathcal{C}(H)$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T$ be the minimal connected induced subgraph of $H$ that spans all critical cliques of $A$, and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$ (see Figure 8). Recall that such a subgraph is unique by Lemma 5.
We first show that $|V(T)|$ contains $O(k^{2})$ critical cliques. By Lemma 5, $|A^{\prime}\setminus A|\leqslant 3\cdot|A|\leqslant 6k$. The connected components of the graph $T\setminus(A\cup A^{\prime})$ are paths since every vertex is of degree at most $2$ and by Lemma 5 there are at most $4k$ such paths. Let $R$ be one of these paths, it is composed of unaffected critical cliques, thus there exists a clean $2$-SC-branch $B$ of $G$ that contains $R$. Moreover, the extremities of $R$ are the attachment points of $B$, hence $B$ have been reduced by Rule 4.5. Thus $R$ is of length at most $k+3$. It remains that $|V(T)|\leqslant 2k+6k+(k+3)\cdot 4k=O(k^{2})$.
We will now show that $\mathcal{C}(H)\setminus V(T)$ contains $O(k^{3})$ critical cliques. First observe that each connected component of $\mathcal{C}(H)\setminus V(T)$ is adjacent to some vertices of $T$ since the graph is reduced by Rule 4.1. Since $\mathcal{C}(H)$ is a block graph, connected components of $\mathcal{C}(H)\setminus V(T)$ are adjacent to either a critical clique of $T$ or a maximal clique of $T$ (else, there would be a diamond in $\mathcal{C}(H)$). We claim that there are $O(k^{3})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to maximal cliques of $T$. Since $T$ is a block graph and $|V(T)|=O(k^{2})$, there are $O(k^{2})$ maximal cliques in $T$. Moreover, there is only one connected component of $\mathcal{C}(H)\setminus V(T)$ adjacent to each maximal clique, otherwise there would be a diamond in $\mathcal{C}(H)$. Take $K$ a maximal clique of $T$ and let $X$ be its adjacent connected component of $\mathcal{C}(H)\setminus V(T)$. Observe that $X$ has to be an union of $1$-SC-branches and their attachment points form a clique. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches in $X$ and each one has been reduced by Rule 4.3, thus $X$ contains at most $24k+$ 2 critical cliques.
Finally, we claim that there are $O(k^{2})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to critical cliques of $T$. First take a critical clique of $T\setminus A$ and its adjacent connected components of $\mathcal{C}(H)\setminus V(T)$, recall that $|T\setminus A|=O(k^{2})$. Observe that they form a $1$-SC-branch, reduced by Rule 4.3, thus the adjacent connected component actually consists in only one critical clique. Next, take a critical clique $a$ of $A$, and let $C_{1},\dots,C_{r}$ be the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to $a$ and $Y$ the union of these connected components (see Figure 8). Observe that each $C_{i}$ has to be an union of $1$-SC-branches and their attachment points form a clique $Q_{i}$. Let $B_{1},\dots,B_{l}$ with attachment points $P_{1},\dots,P_{l}$ be the $1$-SC-branches of $Y$. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches by connected component and if $\Sigma_{i=1}^{l}|P_{i}|>2k+1$, then $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}\leqslant k$, implying that there are at most $13k+$ 1 $1$-SC-branches in $Y$. Each of these $1$-SC-branches is reduced by Rule 4.3, hence $Y$ contains at most $26k+$ 2 critical cliques in total. Since $|A|=O(k)$, there is $O(k)$ such sets of $1$-SC-branches in $\mathcal{C}(H)\setminus V(T)$. Altogether, it remains that $\mathcal{C}(H)\setminus V(T)$ contains $O(k^{2})$ critical cliques adjacent to critical cliques of $T$.
Figure 8: Illustration of the critical clique graph of a reduced instance in the proof of Theorem 4. Square nodes correspond to vertices of $T$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $A^{\prime}\setminus A$.
We conclude that $|V(\mathcal{C}(H))|=O(k^{3})$ and by Lemma 24, we have $|V(\mathcal{C}(G))|\leqslant|V(\mathcal{C}(H))|+2k$, therefore by Rule 4.2 $|V(G)|=O(k^{4})$ holds. To conclude, we claim that a reduced instance can be computed in polynomial time. Indeed, Lemma 14 states that it is possible to reduce exhaustively a graph under Rules 4.1 and 4.2 in linear time. Once this is done, it remains to apply exhaustively Rules 4.3, 4.4 and 4.5 which is ensured by Lemma 23. ∎
5 Kernels for the completion and deletion variants
In this section we present kernels for Strictly Chordal Completion and Strictly Chordal Deletion. It is clear that Rules 4.1 and 4.2 are also safe for these variants of the problem. We can also observe that Rules 4.3, 4.4 and 4.5 are safe for the completion and deletion variants since they are a special case of the edition. However, we can use more efficient rules to manage the $2$-SC-branches and bound their length to a constant number of critical cliques. This is due to only allowing edge addition or edge deletion. In the case of Strictly Chordal Completion, one can observe that for a graph $G$, it is sufficient to consider the clean $2$-SC-branches $B$ of length $3$ which have attachment points that belong to different connected components of $G\setminus V(B^{R})$ since a SC-branch remains connected in any completion of $G$. The next lemma follows from this observation and its proof is contained in the proof of Lemma 21. From this lemma we can devise an improved rule for the completion variant.
Lemma 25.
Let $G=(V,E)$ be graph and $B$ a clean $2$-SC-branch of length at least 3 with attachment points $P_{1},P_{2}$ that are in different connected components of $G\setminus V(B^{R})$. There exists an optimal completion $F$ of $G$ such that:
  • The set of vertices of $B$ affected by $F$ is included in $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$,
  • In $H=G+F$, the vertices of $N_{B}(P_{1})$ (resp. $N_{B}(P_{2})$) are all adjacent to the same vertices of $V(G)\setminus V(B^{R})$.
Rule 5.1.
Let $(G=(V,E),k)$ be an instance of Strictly Chordal Completion and $B$ a clean $2$-SC-branch of $G$ of length at least $3$ with attachment points $P_{1},P_{2}$ that are in different connected components of $G\setminus V(B^{R})$. Remove the vertices $V(B^{R})$ and add two new cliques $K_{1}$ and $K_{2}$ of size respectively $min\{|N_{B}(P_{1})|,k+1\}$ and $min\{|N_{B}(P_{2})|,k+1\}$ with edges $(P_{1}\times K_{1}),(K_{1}\times K_{2}),(K_{2}\times P_{2})$.
In the case of Strictly Chordal Deletion, we can observe that if the two attachment points of $B$, a $2$-SC-branch of length at least $3$, are in the same connected component of $G\setminus V(B^{R})$, then either $B$ is disconnected or the attachment points are not in the same connected component of $(G-F)\setminus V(B^{R})$ for any deletion $F$ of $G$. Indeed, if this was not the case, we could find an induced cycle of length at least $4$ since edge additions are not allowed. The next lemma follows from this observation and its proof is contained in the proof of Lemma 21. As previously we can devise an improved rule from this lemma for the deletion variant.
Lemma 26.
Let $G=(V,E)$ be graph and $B$ a clean $2$-SC-branch of length at least $3$ with attachment points $P_{1},P_{2}$. There exists an optimal deletion $F$ of $G$ such that:
  • If $P_{1}$ and $P_{2}$ are in the same connected component of $(G\triangle F)\setminus V(B^{R})$, then $F$ contains an edge-cut of $B$. This edge-cut is $(P_{1}\times N_{B}(P_{1}))$ or $(P_{2}\times N_{B}(P_{2}))$ or a min-cut of $B$.
  • In each case, the other vertices of $B$ affected by $F$ are included in $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$,
  • In $G-F$ the vertices of $N_{B}(P_{1})$ (resp. $N_{B}(P_{2})$) are all adjacent to the same vertices of $V(G)\setminus V(B^{R})$.
Rule 5.2.
Let $(G=(V,E),k)$ be an instance of Strictly Chordal Deletion and $B$ a clean $2$-SC-branch of $G$ of length at least $3$ with attachment points $P_{1},P_{2}$. Remove the vertices $V(B^{R})$, and add the following path of $5$ cliques:
$K_{min\{|N_{B}(P_{1})|,k+1\}}-K_{k+1}-K_{1}-K_{mc(B)}-K_{min\{|N_{B}(P_{2})|,k+1\}}$
where $K_{n}$ is the clique of size $n$ and $K_{min\{|N_{B}(P_{1})|,k+1\}}$ (resp. $K_{min\{|N_{B}(P_{2})|,k+1\}}$) is adjacent to $P_{1}$ (resp. $P_{2}$).
Lemma 27.
Rules 5.1 and 5.2 are safe and can be applied in polynomial time.
With these new rules, we can observe that in the proof of Theorem 4, the connected minimal induced subgraph $T$ that spans all the affected critical cliques contains $O(k)$ critical cliques. Indeed, the quadratic number of critical cliques in $T$ resulted from the paths in $T$ that contained a linear number of critical cliques. However, with Rules 5.1 and 5.2, these paths contain $O(1)$ critical cliques, implying that $T$ contains $O(k)$ critical cliques. With this observation, it follows that a reduced graph contains $O(k^{2})$ critical cliques and thus contains $O(k^{3})$ vertices.
Theorem 5.
Strictly Chordal Completion and Strictly Chordal Deletion admit kernels with $O(k^{3})$ vertices.
6 Conclusion
We presented polynomial kernels for the edition and deletion variants of block graph edge modification problems and the three variants of strictly chordal edge modification problems. Our conviction is that the approach based on decompositions of the target class, combined with the ability of reducing the size of the bags of the decomposition and of limiting the number of affected bags to $O(k)$ is a promising starting point for edge modification problems, especially into subclasses of chordal graphs. The technique has been employed especially for classes that admit a tree-like decompositions with disjoint bags (e.g., 3-leaf power [4], trivially perfect graphs [21, 19] and $\mathcal{M}$-free graphs [30]), also for other types of tree-like decompositions with non-disjoint bags (e.g., ptolemaic graphs [14]). We generalize it here to strictly chordal graphs, that have a decomposition into disjoint bags as nodes of a block graph.
The difficulty is that, at this stage, each class requires ad-hoc arguments and reduction rules, based on its specific decomposition. An ambitious goal would be to obtain a generic algorithm for edge modification problems into any class of chordal graphs, plus a finite set of obstructions, as conjectured by Bessy and Perez [5] for completion problems. As an intermediate step, we ask whether $4$-leaf power completion, deletion and editing problems admit a polynomial kernel.